OpenStudy (anonymous):

find the following limits, or state that they do not exist. a) \[\lim_{y \rightarrow 2}\frac{ y^4-16 }{ y^4+2y^3-y^2-2y }\] b)\[\lim_{t \rightarrow -2^+}\sqrt[4]{8+t^3}\] c) \[\lim_{x \rightarrow 1^+}\left| x-1 \right|/ x-1\]

ganeshie8 (ganeshie8):

factor and see if any bad stuff cancels out

ganeshie8 (ganeshie8):

actually there are no bad factors, you can directly plugin y=2 into the expression and evaluate

OpenStudy (anonymous):

\[\lim_{y \rightarrow 2} (y-2)(y+2)(y^2+4) / y(y^3+2y^2-y-2) \] I'm stuck here.

OpenStudy (dan815):

the bad stuff lol..

OpenStudy (anonymous):

oh, there's no bad stuff... how you know that?

ganeshie8 (ganeshie8):

good question :) you will get to know only after plugging in y=2 it is "bad" if you get 0 on both numerator and denominator

ganeshie8 (ganeshie8):

because \(\large \dfrac{0}{0}\) is one of the indetermnate forms..

OpenStudy (anonymous):

oh yeah! forgot about that. So for a) is 0?

ganeshie8 (ganeshie8):

looks good!

OpenStudy (anonymous):

is b) also 0? but why is there t -->-2^+?

ganeshie8 (ganeshie8):

that means you're walking on x axis from right of -2

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

\[\lim_{t \rightarrow -2^+}\sqrt[4]{8+t^3} =\sqrt[4]{\lim_{t \rightarrow -2^+}8+t^3} \]

OpenStudy (anonymous):

love your drawing :)

ganeshie8 (ganeshie8):

plugin t = -2 and evaluate \[8 + (-2)^3 = 8-8 = 0\] so we get \[\lim_{t \rightarrow -2^+}\sqrt[4]{8+t^3} =\sqrt[4]{\lim_{t \rightarrow -2^+}8+t^3} = \sqrt[4]{0} = 0\]

ganeshie8 (ganeshie8):

<3

OpenStudy (anonymous):

oh i see..

OpenStudy (anonymous):

I'm sure how to do c) Forgot the definition of absolute value.

ganeshie8 (ganeshie8):

Easy.. absolute value can never be negative, so we define it in the best way we can think of : |x| = x if x >= 0 |x| = -x if x < 0

ganeshie8 (ganeshie8):

tell me this is `x-1` positive when x is greater than 1 ?

OpenStudy (anonymous):

yes